# The Language of Interoception: Examining Embodiment and Emotion Through a Corpus of Body Part Mentions

### v1

@@ 5 Research Questions on BPMs–Affect
The primary goal of our work is to explore how language can shed light on the connection between the body, emotion, and well-being. In this section we explore how BPMs are associated with emotions. We explore this question using emotions associated with words that co-occur with BPMs (using large word–emotion association lexicons) as well as perceived emotions of the speaker (using the new human-annotated Spinn3rBPM-Zhuang dataset we introduced earlier).
@@
Method: This question aims to shed light on whether the relationship between emotion and embodiment manifests in social media text. We use large word–emotion association lexicons to look at whether words that are associated with particular emotions co-occur with body part words. We use this method both because it is an interpretable approach to understanding the emotional associations of BPMs, and because it is a computationally inexpensive method which has been shown to effectively capture aggregate emotion arcs (changes in emotions) across large samples Teodorescu and Mohammad (2023). In this experiment, we look at the proportion of samples (tweets/blog posts) containing at least one word associated with various emotion categories: anger, anticipation, disgust, fear, joy, sadness, surprise, and trust from Plutchik’s set of emotions Plutchik (2001), and high or low valence (positive–negative), arousal (calm–sluggish), and dominance (in control–out of control). We obtain the word–emotion associations from the NRC Emotion Lexicon Mohammad and Turney (2013, 2010) and the NRC VAD Lexicon Mohammad (2018). We compute these proportions for the “my <BPM>", "his/her/their <BPM>", "your <BPM>", as well as the no BPM corpora.
@@
Results: Figure 4 shows the results for valence, arousal and dominance. We find that instances containing BPMs have markedly higher percentages of emotion-associated co-occurring words than instances not containing BPMs. This tends to be true for all corpus–emotion dimension pairs except for TUSC–High Dominance. We find the same pattern for categorical emotions (Figure 12 in the Appendix), except for TUSC–Trust, where “my <BPM>" instances have slightly lower scores than no BPM instances. We also observe that the jump in scores from the no BPM corpus to BPM corpora is highest for the low-valence and low-dominance dimensions ($\sim$15 percentage points).
@@
Method: Body parts are often referenced as physically involved in emotional responses (e.g., my heart skipped a beat, my stomach dropped). In BA1, we showed that BPM instances are more likely to have emotion-associated co-terms than their no-BPM counterparts. In this question, we explore whether this increase only exists in explicitly embodied emotion. To do this, we analyze samples which are human-annotated by Zhuang et al. (2024) as either containing embodied emotion (where a body part physically participates in expressing the emotion—annotated as embodied) or not (annotated as non-embodied). Specifically, we looked at the degree of emotion-word co-occurrences in the embodied samples and in the non-embodied samples. As a separate and complementary experiment, we manually annotated these exact instances for whether the speaker was feeling any of the emotion categories (anger, fear, joy, sadness, surprise, and trust), by crowdsourcing on Amazon Mechanical Turk. This allows us to determine, in an arguably a more direct way than lexical co-occurrences, whether embodied BPM instances tended to be more emotional than non-embodied BPM instances.
Results: Figure 5 shows the percentages of embodied and non-embodied samples where the speaker is experiencing an emotion. Figure 6 shows the percentages of embodied and non-embodied samples that include at least one word associated with an emotion." We find that both methods indicate no notable difference in the percentages of emotional samples across explicitly embodied versus non-embodied (or more precisely, not explicitly embodied) BPM instances (<4 percentage points). In contrast, there is a stark difference between the the emotion percentages of the no BPM samples and the embodied/non-embodied BPM samples ) (>30 percentage points).
Discussion: The results show that emotional words appear more frequently in BPM sentences, regardless of whether or not the BPMs are physically linked to the emotion (explicitly embodied). Speakers are also equally likely to be expressing emotion whether the BPM is embodied or not. Overall, this supports a deeper connection between BPMs and affect, independent of whether there is an explicit description of a physical role of of the BPM in the emotion.
@@
Method: In BA1, we looked at the co-occurrence of posts containing “my <BPM>" with emotion-associated words. Here, we are interested in comparing this average score to posts containing specific “my <BPM>" types. For the most common “my <BPM>" types that pass a threshold of minimum appearance of 100 instances across each corpus, we calculated the average proportion of samples that include words associated with specific emotion dimensions. We calculated a mean and standard deviation over these proportions and use these values to find “my <BPM>" types which are significantly associated with particular emotional dimensions. We also calculated the standard deviation for each emotional category across all common “my <BPM>" types to identify which types which are particularly associated with certain emotional dimensions.
@@
Results: We find that different “my <BPM>" types are associated with different emotions to markedly different degrees, and that different profiles of associations for the same type can be found in different corpora. However, some “my <BPM>" types carry consistent cross-corpus associations, such as "my stomach" being most associated with sadness in both TUSCctry-BPM and Spinn3rBPM, whereas "my chest" is most associated with anger. (Proportions for emotional word co-occurrence across “my <BPM>" types for TUSCctry-BPM and Spinn3rBPM are shown in Figures 13 through 16 and the most associated emotion for each of the BPMs—which are often negative—are shown in Table 10 in the Appendix.)
Discussion: These results indicate that referral to one’s body parts are associated with different affective expressions online. The overall trends in $TUSC_{city}$ seem to imply that referral to one’s own body parts online often arise from situations of pain, lethargy, and a lack of control.
@@ 6 Do BPMs Correlate with Health?
Method: To evaluate this hypothesis, we look at available city-wide health data City-Health-Dashboard-Dataset (2025) for all 25 American cities in the TUSCcity dataset, and correlations between the proportion of regional tweets containing “my <BPM>"/“BPM" and four health measures: frequent mental distress, frequent physical distress, life expectancy, and physical inactivity. [^3] https://www.cityhealthdashboard.com As points of baseline comparison, we also look at how the health factors investigated are correlated with the number of tweets from each region and the correlation between the proportion of emotion-associated words (from the NRC Emotion/VAD lexicons) with the health outcomes.
Results: Table 3 shows the Spearman rank correlations as well the p-values (we consider the correlations to be statistically significant if the p-value is below 0.05). Observe that the number of tweets per city is not correlated with the health outcomes (row a). We find that most emotion–health outcome pairs are also not correlated or only slightly correlated. The highest correlation numbers are for fear–physical activity (see row b). (Table 11 in the Appendix shows correlations for each of the emotion–health outcome pairs.) In contrast, the proportion of “<BPM>" and “my <BPM>" mentions (rows c and d) are moderately or strongly correlated with all three negative health outcomes and anticorrelated with life expectancy (the statistically significant results are shown in bold). Notably, frequent physical distress and physical inactivity are remarkably correlated with higher myBPM usage (Spearman’s $r$ = 0.721, Spearman’s $r$ = 0.704 respectively), and life expectancy is strongly negatively correlated with BPM use (Spearman’s $r$ = -0.613). Overall, these results show that simple metrics capturing the proportion of mentions of body parts in social media can be useful indicators of both physical and mental health.
@@ 7 Conclusion
We created novel corpora designed specifically for the study of Body Part Mentions (BPMs), which includes the first-ever dataset of samples explicitly annotated for the emotions of human entities possessing BPMs. Using these corpora, we answered a series of research questions on the significance of body-related words in everyday language, the relationship between embodiment and emotion, and factors correlated with BPM frequency such as emotional context, time of week/year, and region. We showed that BPMs occur frequently in social media texts and have notable temporal and geographic trends. We also showed that BPM instances have markedly higher emotion associations than non-BPM instances—with an especially marked increase in low valence (negativity) and low dominance (helplessness) instances. Most notably, through experiments on data from 25 US cities, we showed that the degree of BPM usage is a powerful indicator of aggregate-level well-being. Although the connection between language, embodiment, and affect is now well-established, this paper is – to our knowledge – the first-ever approach to understanding this relationship grounded in large amounts of language data. We release our BPM corpus to the public, and hope that our work demonstrates body-related language as a rich and interesting source of material for future NLP research to investigate the deeper connection between language, embodiment, and emotional wellbeing.
@@ Acknowledgments
Many thanks to Tara Small for helpful discussions and comments.
@@ Appendix A Full Dataset Descriptions
All relevant datasets and subsets to this paper can be viewed in Table 4.
@@ Appendix B Obtaining human ratings of emotion for Spinn3rBPM-Zhuang corpus
For each instance, we identify all possible BPMs. For each sample presented, we ask the crowdworker to identify whether a BPM in the sample belongs to the "speaker" or the "non-speaker". We also present them a description of the emotions we will want annotations for (See Figure 7). For example, in the sentence "Robin placed her hand on Kevin’s shoulder", we would tell the annotator to identify the owner of "hand" or "shoulder", which would lead them to annotate for the emotion of Robin or Kevin respectively. We also note that all samples from the Spinn3rBPM-Zhuang include BPMs that are preceded by the possessive pronoun "my", "his", or "her", guaranteeing that there is always an entity whose emotional state can be inferred from the BPM.
We then present six emotional categories that they can annotate from, along with descriptions of these emotional categories (Figure 8). For each emotional category, they are five ranked categories they can choose from to indicate the severity of the emotion (no/slight/moderate/high/very high) (Figure 9).
@@ Appendix C Most Frequent myBPMs by Corpus (Supplementary Table, B3)
The top 20 most frequent myBPMs in each corpus, along with their frequency relative to all myBPMs present in their respective corpus, can be viewed in Table 5.
@@
Appendix D BPM scores by country (Supplementary Figures, B5)
Figure 10 shows the percentage of tweets containing at least one instance of “my <BPM>" from Canada and the USA between 2015 and 2021.
Figure 11 displays the monthly percentage of tweets with at least one “my <BPM>" from Canada and the USA between 2015 and 2021.
@@ Appendix E Are mentions of the body associated with longer utterances: longer sentences/tweets?
[…]
@@ Appendix F Emotion associations between myBPMs, yourBPMs, 3pBPMs, noBPMs (Supplementary Figures, BA1)
Figure 12 shows the percentage of sentences containing at least one positive word associated with Plutchik’s eight emotional categories (NRC emotion lexicon) across the same BPM categories in each corpus.
@@
Appendix G Controling for Post Length (Supplementary Tables, BA2)
@@
Note that tweets are limited to 280 characters. However, blog posts can be longer. Tables 8 and 9 in the Appendix show the percentage of emotion-associated co-occurring words when controlling for blog post length for VAD and emotion categories, respectively.

### v2

@@ 5 Research Questions on BPMs–Affect
The primary goal of our work is to explore how language can shed light on the connection between the body, emotion, and well-being. In this section we explore how BPMs are associated with emotions. We explore this question using emotions associated with words that co-occur with BPMs (using large word–emotion association lexicons) as well as perceived emotions of the speaker (using the new human-annotated Spinn3rBPM-Zhuang dataset we introduced earlier). In this section, we take a special interest in samples including "my <BPM>", since our emphasis is rooted in theories of embodied emotion and health psychology, which suggest that references to one’s own body are more likely to reflect internal emotional and physical states.
@@
Method: This question aims to shed light on whether the relationship between emotion and embodiment manifests in social media text. In this experiment, we look at the proportion of samples (tweets/blog posts) containing at least one word associated with various emotion categories: anger, anticipation, disgust, fear, joy, sadness, surprise, and trust from Plutchik’s set of emotions Plutchik (2001), and high or low valence (positive–negative), arousal (calm–sluggish), and dominance (in control–out of control). We obtain the word–emotion associations from the NRC Emotion Lexicon Mohammad and Turney (2013, 2010) and the NRC VAD Lexicon Mohammad (2018b). We compute these proportions for various BPM categories: “my <BPM>”, "his/her/their <BPM>", "your <BPM>", as well as the no BPM corpora.
@@
Results: Figure˜5 shows the results for valence, arousal and dominance. We find that instances containing BPMs have higher percentages of emotion-associated co-occurring words than instances not containing BPMs. This is true across 36/42 corpus–dimension pairs. We also observe that the jump in scores from the no BPM corpus to BPM corpora is highest for the low-valence and low-dominance dimensions ($\sim$15 percentage points). We find that the BPM category has a statistically significant effect on the percentage of samples containing emotion-associated terms (more details on these tests can be found in Appendix˜I).
@@
Method: Body parts are often referenced as physically involved in emotional responses (e.g., my heart skipped a beat, my stomach dropped). Prior work (i.e. Zhuang et al.) has assumed that such samples in natural language may be responsible for the emotional associations between emotional and body-related language. In BA1, we showed that BPM instances are more likely to have emotion-associated co-terms than their no-BPM counterparts. With this question we explore whether this increase only exists in explicitly embodied emotion. To do this, we analyze samples which are human-annotated by Zhuang et al. (2024) as either containing embodied emotion (where a body part physically participates in expressing the emotion—annotated as embodied) or not (annotated as non-embodied). Specifically, we looked at the degree of emotion-word co-occurrences in the embodied samples and in the non-embodied samples. As a separate and complementary experiment, we manually annotated these exact instances for whether the speaker was feeling any of the emotion categories (anger, fear, joy, sadness, surprise, and trust), by crowdsourcing on Amazon Mechanical Turk. We note a high inter-annotator agreement rate on our human annotations using Split-Half Class Match Percentage (90.1% on 2 bins to 69.2% on 10 bins). Further details on annotator agreement can be found in Appendix˜B. This allows us to determine the extent to which explicitly embodied BPMs are more or less emotional than not explicitly embodied BPM instances (non-embodied for short).
Results: Figure˜6 shows the percentages of embodied and non-embodied samples where the speaker is experiencing an emotion. Figure˜7 shows the percentages of embodied and non-embodied samples that include at least one word associated with an emotion. We find that both methods indicate no notable difference in the percentages of emotional samples across explicitly embodied versus non-embodied (or, more precisely, not explicitly embodied) BPM instances (<4 percentage points). In contrast, there is a stark difference between the the emotion percentages of the no BPM samples and the embodied/non-embodied BPM samples ) (>30 percentage points).
Discussion: The results show that emotional words appear more frequently in BPM sentences, regardless of whether or not the BPMs are explicitly written as physically connected to emotion in in the text (explicitly embodied). Speakers are also equally likely to be expressing emotion whether the BPM is embodied or not. This unexpected finding supports the theory of constructed emotion and highlights a stronger connection between emotional expression and body-related language than theorized by previous works.
@@
Method: In BA1, we looked at the co-occurrence of posts containing “my <BPM>” with emotion-associated words. Here, we are interested in comparing this average score to posts containing specific “my <BPM>” types. For the most common “my <BPM>” types that occur in at least 100 instances across each corpus, we calculated the average proportion of instances that include words associated with specific emotion dimensions. We calculated a mean and standard deviation over these proportions and use these values to find “my <BPM>” types which are significantly associated with particular emotional dimensions. We also calculated the standard deviation for each emotional category across all common “my <BPM>” types to identify which types are markedly associated with certain emotional dimensions.
@@
Results: We find that different “my <BPM>” types are associated with different emotions to markedly different degrees, and that different profiles of associations for the same type can be found in different corpora. However, some “my <BPM>” types carry consistent cross-corpus associations, such as my stomach being most associated with sadness in both TUSCctry-BPM and Spinn3rBPM, whereas my chest is most associated with anger. (Proportions for emotional word co-occurrence across “my <BPM>” types for TUSCctry-BPM and Spinn3rBPM are shown in Figures 13 through 16 and the most associated emotion for each of the BPMs—which are often negative—are shown in Table˜8 in the Appendix.)
Discussion: These results indicate that referral to one’s body parts are associated with different affective expressions online. Overall trends in $TUSC_{city}$ seem to imply that referral to one’s own body parts online often arise from situations of pain, lethargy, and a lack of control.
@@ 6 Do BPMs Correlate with Health?
Method: To evaluate this hypothesis, we look at available city-wide health data City-Health-Dashboard-Dataset (2025) for all 25 American cities in the TUSCcity dataset, and correlations between the proportion of regional tweets containing “my <BPM>”/“BPM" and four health measures: frequent mental distress, frequent physical distress, life expectancy, and physical inactivity. [^8] https://www.cityhealthdashboard.com As points of baseline comparison, we also look at how the health factors investigated are correlated with the number of tweets from each region, and the correlation between the proportion of emotion-associated words (from the NRC Emotion/VAD lexicons) with the health outcomes studied.
Results: Table˜3 shows the Spearman rank correlations as well the p-values (we consider the correlations to be statistically significant if the p-value is below 0.05). Observe that the number of tweets per city is not correlated with the health outcomes (See Row A). We find that most emotion–health outcome pairs are also not correlated or only slightly correlated. The highest correlation numbers are for fear–physical activity (See Row B). (Table˜9 in the Appendix shows correlations for each of the emotion–health outcome pairs.) In contrast, the proportion of “<BPM>" and “my <BPM>” mentions (rows c and d) are moderately or strongly correlated with all three negative health outcomes and anticorrelated with life expectancy (statistically significant correlations are bolded). Notably, frequent physical distress and physical inactivity are remarkably correlated with higher myBPM usage (Spearman’s $r$ = 0.721, Spearman’s $r$ = 0.704 respectively), and life expectancy is strongly negatively correlated with BPM use (Spearman’s $r$ = -0.613). Overall, these results show that simple metrics capturing the proportion of mentions of body parts in social media can be useful indicators of both physical and mental health. While our findings reveal a statistically significant correlation between body part mentions (BPMs) and regional health outcomes, we do not claim a direct causal relationship. It is likely that both language use and health indicators are shaped by broader social and demographic factors—such as educational attainment, economic status, or regional linguistic norms—which may contribute to the observed patterns. We see this possibility as an exciting result – rather than treating BPMs as independent predictors of health, we interpret these correlations as evidence of shared variance that may offer insight into the sociolinguistic embedding of embodied experience. We highlight these associations as a starting point for future research that more directly models such confounding factors.
@@ 7 Conclusion
We created novel corpora designed specifically for the study of Body Part Mentions (BPMs), which includes the first-ever dataset of samples explicitly annotated for the emotions of human entities possessing BPMs. Using these corpora, we answered a series of research questions on the significance of body-related words in everyday language, the relationship between embodiment and emotion, and factors correlated with BPM frequency such as emotional context, time of week/year, and region. We showed that BPMs occur frequently in social media texts and have notable temporal and geographic trends. We also showed that BPM instances have markedly higher emotion associations than non-BPM instances—with an especially marked increase in low valence (negativity) and low dominance (helplessness) instances. Most notably, through experiments on data from 25 US cities, we showed that the degree of BPM usage can is a powerful indicator of aggregate-level well-being. Although the connection between language, embodiment, and affect is now well-established, this paper is – to our knowledge – the first-ever approach to understanding this relationship grounded in large amounts of language data. We release our BPM corpus to the public, and hope that our work demonstrates body-related language as a rich and interesting source of material for future NLP research to investigate the deeper connection between language, embodiment, and emotional wellbeing. Although BPMs are a relatively simple tool for investigating the relationship between embodiment and everyday language, they offer a scalable, interpretable signal for understanding this connection empirically. We hope that our initial exploratory work, through showing that body-related language carries diverse and meaningful associations, emphasizes the richness of studying the intersection of embodiment and natural language.
@@
Since the primary aim of this paper is descriptive rather than explanatory to highlight that there are diverse and meaningful associations which BPMs carry, warranting future study, we do not isolate any specific mechanisms (i.e. effects of social variables, classification of BPM usage types, or deeper linguistic analysis) which could explain these associations. We encourage future work which can further probe the exact causes and explanations for the associations which we have discovered in this paper.
For this reason, we chose to use lexicon-based approaches to studying affect in our corpus, since although they cannot capture the full nuance of emotional expression in a single sample, especially in figurative or context-dependent language, they remain valuable tools for capturing broad trends across large corpora, and their interpretability makes them suitable for an exploratory, large-scale study like ours. such as ML-based classifiers, could possibly help provide more fine-grained and context-sensitive emotion detection for more specific and contextual research questions (compared to the more general and exploratory research questions we attempt to answer in this paper).
@@ Acknowledgments
Many thanks to Tara Small for helpful discussions and comments. We also thank Lisa Pennel for her feedback and support on this paper.
@@ Appendix A Full Dataset Descriptions
All relevant datasets and subsets to this paper can be viewed in Table˜4.
@@ Appendix B Obtaining human ratings of emotion for Spinn3rBPM-Zhuang corpus
For each instance, we identify all possible BPMs. For each sample presented, we ask the crowdworker to identify whether a BPM in the sample belongs to the "speaker" or the "non-speaker". We also present them a description of the emotions we will want annotations for (See Figure˜8). For example, in the sentence "Robin placed her hand on Kevin’s shoulder", we would tell the annotator to identify the owner of "hand" or "shoulder", which would lead them to annotate for the emotion of Robin or Kevin respectively. We also note that all samples from the Spinn3rBPM-Zhuang include BPMs that are preceded by the possessive pronoun "my", "his", or "her", guaranteeing that there is always an entity whose emotional state can be inferred from the BPM.
We then present six emotional categories that they can annotate from, along with descriptions of these emotional categories (Figure˜9). For each emotional category, they are five ranked categories they can choose from to indicate the severity of the emotion (no/slight/moderate/high/very high) (Figure˜10).
@@
We assess annotation reliability using the Split-Half Class Match Percentage (SHCMP) as reported in the paper, a method adapted from traditional split-half reliability to handle categorical labels like those used for emotion intensity. SHCMP evaluates how consistently items are classified across multiple random groupings of the dataset. Specifically, the data is divided into n random subsets (with n = 2 representing a typical half-split) 1,000 times, and the average proportion of items that receive the same label across these splits is computed. A higher SHCMP score reflects greater reliability, indicating that the labels are likely to remain stable across repeated annotations.
@@ Appendix C Most Frequent myBPMs by Corpus (Supplementary Table, B3)
The top 20 most frequent myBPMs in each corpus, along with their frequency relative to all myBPMs present in their respective corpus, can be viewed in Table˜5.
@@
Appendix D BPM scores by region (Supplementary Figures and Statistical Tests, B5)
Figure 11 displays the monthly percentage of tweets with at least one “my <BPM>” from Canada and the USA between 2015 and 2021.
To assess whether the city of a post influences the probability of including a body-part mention (BPM), we modeled the number of posts containing BPMs out of the total posts per city using a binomial logistic regression, with city included as a categorical predictor. This approach models the log-odds of a post containing a BPM as a function of city. We then performed a likelihood ratio test comparing this model to a null model containing only an intercept, which indicated that cities overall have a highly significant effect on BPM usage (LR = 357.61, p < $10^{-67}$). When performing the same test using country as a categorical predictor, we find statistically significant results as well (LR = 9.8, p = 0.002), although its effect is considerably weaker than that of individual cities.
@@ Appendix E Are mentions of the body associated with longer utterances: longer sentences/tweets?
Results: In Spinn3r, sentences containing BPMs are substantially longer than sentences without BPMs, with an average length of 482.07 characters compared to 146.53 characters for non-BPM sentences, making the average BPM sentence length 3.29x the length of the average non-BPM sentence. In the TUSC datasets, a similar trend holds but with a smaller magnitude: BPM tweets average 130.99 characters in TUSCcity{city} (vs. 93.98 for non-BPM) and 133.42 characters in TUSCctry{ctry} (vs. 101.45 for non-BPM), corresponding to a 1.39x and 1.32x increase, respectively.
@@ Appendix F Emotion associations between myBPMs, yourBPMs, 3pBPMs, noBPMs (Supplementary Figures, BA1)
Figure˜12 shows the percentage of sentences containing at least one positive word associated with Plutchik’s eight emotional categories (NRC Emotion Lexicon) across the same BPM categories in each corpus.
@@
Appendix G Controlling for Post Length (Supplementary Tables, BA2)
@@
Note that tweets (which we use as individual posts for our tweet dataset, TUSC) are limited to 280 characters. However, blog sentences (which we use as individual posts for our blog dataset, Spinn3r) can be longer. Tables 6 and 7 in the Appendix show the percentage of samples co-occuring with emotion-associated words when controlling for blog post length for VAD and emotion categories, respectively.
Appendix H Effect of time on BPM usage (Supplementary statistical tests, B4)
To assess the statistical significance of seasonal variation in BPM usage, we modeled the log-odds of a post containing a BPM as a function of month using a binomial logistic regression. We treat month as a continuous variable, starting from April, since this is the month where we can start to observe the trend of decreasing BPM use throughout the year. The results indicate a statistically significant decline in BPM usage over the months following April (coefficient = -0.0036, p < 0.001). When performing the same test with weekday, starting from Monday, we find a statistically significant decline in BPM usage from Monday as well (coefficient = -0.0104, p < 0.001).